Shift report 20011014 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 14 2001 - 23:51:36 EDT

  • Next message: Apache: "Shift report 20011015 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Oct 14 23:51:35 2001, Shift: 16:00-24:00
    Supervisor: CEJ, with: DO
    Log book#10 pages 7 - 9
    
    Shift Summary:
    ------------------------------------------------------------
    1. Nice long store approx 5 hours from 17:00. Beam lost
       without warning at 22:03. Quench link interlock.
    
    2. Data runs:
       5338  ( 249K trigger 1 )
       5339  ( 212K trigger 1 )
       5340  ( 198K trigger 1 )
       5341  (  44K trigger 1 )
       Total : 703K trigger 1 
     
       Pedestal runs: 5334, 5341 
       TPC pedestal runs: 5335, 5342
    
    3. Turned up voltages on T1 Fibers and T2 Fibers. Checked
       the ADC signal on the scope and it looks nice! The data
       is available from run 5338. T3 anode1, T4 anode1 and
       T5 anode3xy tripped as the beam was lost. 
        
    
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 14 2001 - 23:52:10 EDT